1)Calculate K at 25 C for the reaction below (i.e. in question 7).

Correct answer is 5.6E27 +/- 0.3E27, but I am unable to figure out how to get it.

2) Calculate ΔGo for the reaction below (questions 7)

------------------------------------------------------------------------------------------------------------------------------------------------

Question 7

Under standard conditions, what product(s) form when Cl2 gas is bubbled into a solution of NaI

a) no reaction
b) I2 (Correct Answer)
c) Na+
d) Na

Cl2 + 2e- -> 2Cl 1.358V

2I -> I2 + 2e -0.535V

1.358V - 0.535 V = 0.823V (Use this to solve question asked)

Since E Cell is 0.823V (positive) I2 will be produced.
Answer #1

The reaction involved is

Cl_{2}\: + 2NaI \: \rightarrow I_{2} \: +2NaCl

so iodine gas is released.

given Ecell = + 0.893V

n= number of electrons involved = 2

T = 25oC = 323K

then

\Delta G = -nFE_{cell}

\Delta G = -2X96500X0.893 =\: - 172349joules =\: -172.34kj

and K is given by

log K = \: - \frac{\Delta G}{2.303RT}

log K = \: - \frac{\left ( -172349 \right )}{2.303X8.314X 298}= +30.2

k = antilog 30.2 = 1.584X10^{30}